# and bases of the Temperley–Lieb algebra

### v1

@@ Abstract.
VP was partially supported by the Spanish project PID2022-137283NB-C21 of MCIN/AEI/10.13039/501100011033 / FEDER, UE, by the Spanish–German project COMPOTE (AEI PCI2024-155081-2 & DFG 541393733), by the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M), by the Departament de Recerca i Universitats de la Generalitat de Catalunya (2021 SGR 00697), and by the French–Austrian project PAGCAP (ANR-21-CE48-0020 & FWF I 5788).
@@ 1. Introduction
Our work starts from the observation that this excedance relation on the Bruhat order actually extends to a congruence of the lattice of alternating sign matrices (which is the MacNeille completion of the Bruhat order). The congruences of a lattice are the equivalence relations which respect the lattice structure, and they come with rich combinatorial and algebraic properties. Our main results (partially) extend the results of [BG24] to all lattice congruences $\equiv$ on alternating sign matrices whose quotient is isomorphic to the Stanley lattice, which we call catalan congruences.
@@
A few comments on this statement. In (i), the covexillary permutation matrices that appear as maxima of congruence classes depend on the congruence. However, any covexillary permutation matrix appears as the maxima of a class for at least one catalan congruence. In (ii), the minimal alternating sign matrix of each class is not necessarily a permutation (we actually discuss along the paper the specific congruences with this property). But the permutations in each congruence class form an interval of the Bruhat order, whose minimum is $321$-avoiding and whose maximum is covexillary, and (by counting) all $321$ permutations appear as minimal permutations of congruence classes. Finally, we note that the problem to extend Theorem 1.1 (iv) to any catalan congruence will be investigated in a future paper.
Our approach to Theorem 1.2 is a combination of lattice and geometric perspectives on alternating sign matrices. It is well known that the lattice of alternating sign matrices is distributive, hence isomorphic to the inclusion lattice of lower sets of its join irreducible subposet by the fundamental theorem for distributive lattices. As already observed and exploited in [Pro01, Str11], the join irreducible poset of the lattice of alternating sign matrices can be seen geometrically as a poset on the integer points inside the (n2)(n-2)th dilate of a $3$-dimensional tetrahedron resting on an edge (see Figure 3). Pushing slightly further this geometric perspective in Section 2.3, we interpret a lower set in this tetrahedral poset as a stack of rhombic tetrahedra (see Figure 8). The upper hull of this stack is the colored polygonal surface separating the lower set from its complementary upper set (see Figure 5). Projecting this surface to the ground (see Figure 6) naturally leads to interpret geometrically the classical combinatorial models in bijection with alternating sign matrices (e.g. height functions, six-vertex model, fully packed loop), and to cook up some seemingly new models (e.g. cliff model). We note that this perspective also naturally leads to higher dimensional analogues of alternating matrices which will deserve further investigation.
@@ 2.2. Dyck paths
A Dyck path of semilength $n$ is a path with up steps $(1,1)$ and down steps (1,1)(1,-1), starting at $(0,0)$, ending at $(n,n)$, and never passing strictly below the horizontal axis. They define the Dyck path lattice $\mathbf{Dyck}_{n}$ (also known as the Stanley lattice) with the relation $P\leq Q$ if $P$ stays below $Q$. See Figure 2 (left). The join (resp. meet) irreducible Dyck paths are those with a single peak (resp. valley), hence the poset $\mathcal{J}(\mathbf{Dyck}_{n})$ (resp. $\mathcal{M}(\mathbf{Dyck}_{n})$) is isomorphic to the triangular poset $\mathcal{D}_{n}$ of triples $(y_{1},y_{2},y_{3})$ with $y_{1}+y_{2}+y_{3}=n-2$, ordered by $(y_{1},y_{2},y_{3})\leq(z_{1},z_{2},z_{3})$ if and only if $y_{1}+y_{2}\leq z_{1}+z_{2}$ and $y_{2}+y_{3}\leq z_{2}+z_{3}$. See Figure 3 (left). (Note that this poset could be encoded using two coordinates. However, we prefer to work with barycentric coordinates throughout the paper, as they make the presentation more symmetric.)
@@ 2.3. Alternating sign matrices
Note that via the above-mentioned bijection between ASMs and HFMs, we can transport the lattice $\mathbf{HFM}_{n}$ on HFMs of order $n$ to a lattice $\mathbf{ASM}_{n}$ on ASMs of order $n$. See Figure 2 (right). This lattice is the MacNeille completion of the Bruhat order on permutations of $[n]$. We call ascents and descents of an ASM the ascents and descents of its corresponding HFM. The join (resp. meet) irreducible ASMs are known to be the matrices of the bigrassmannian (resp. anti-bigrassmannian) permutations, i.e. the permutations $\sigma$ such that $\sigma$ and its inverse $\sigma^{-1}$ have only one descent (resp. ascent). Namely, an integer point $\boldsymbol{x}=(x_{1},x_{2},x_{3},x_{4})$ in the tetrahedron corresponds to the bigrassmannian and anti-bigrassmannian permutations
@@
It is also interesting to represent an ASM as the surface $S$ separating the lower set $X$ of $\mathcal{T}_{n}$ and its complement. For this, we include additional edges so that each integer point of the tetrahedron gets $12$ incident arrows (one in the direction $\boldsymbol{e}_{i}-\boldsymbol{e}_{j}$ for each $i\neq j\in[4]$). The horizontal edges, in direction $\pm(1,0,0,-1)$ and $\pm(0,1,-1,0)$, are colored in gray. The additional endpoints are considered in $X$ (resp. in its complement) if they are incident to an arrow which is also incident to one of the two bottom (resp. top) faces of the tetrahedron. We then add a rhombus orthogonal to each arrow from a point in $X$ to a point in its complement (the color of the rhombus matches the color of the orthogonal arrow). The corresponding surface is illustrated in Figure 5 (middle and right).
@@
    Note that, up to recoloring, this rainbow tiling is equivalent to a 4-coloring of the aztec diamond.
@@
  • The six-vertex model is given by the horizontal diagonals of the rhombi of $S$, oriented counterclockwise around the surface $S$ (in other words, in the direction of the march of the dahu [^1] https://en.wikipedia.org/wiki/Dahu), and seen from above. See Figure 6 (c). Equivalently, it can be obtained by placing south/west/north/east pointing arrows in the red/blue/orange/green tiles of the rainbow tiling of Figure 6 (a). This gives an orientation of the integer grid such that
@@
    It is called six-vertex model as it has $6$ types of vertices: [Uncaptioned image]. Note that, as the nonzero entries of the ASM correspond to saddle points of $S$, they correspond to the vertices of the six-vertex model with opposite incoming arrows and opposite outgoing arrows. See the connection between Figure 6 (b) & (c).
  • The six-square model is given by the non-horizontal diagonals of the rhombi of $S$, oriented upwards, and seen from above. See Figure 6 (d). Equivalently, it can be obtained by placing east/south/west/north pointing arrows in the red/blue/orange/green tiles of the rainbow tiling of Figure 6 (a). Equivalently, it is the dual directed graph of the six-vertex model of Figure 6 (c) (meaning that each arrow is rotated by $90^{\circ}$). This gives an orientation of the half-integer grid such that
@@
    We call it six-square model as there are $6$ types of squares: [Uncaptioned image].
  • The peak–pit matrix is given by the local optima of the surface $S$, seen from above. Figure 6 (e). They can be visualized as the rainbow half integer points of the rainbow tiling of Figure 6 (a). Equivalently, they are the points with indegree or outdegree $4$ in the six-square model of Figure 6 (d). We are not aware that this model has been studied before.
  • The six-vertex configuration is given by the horizontal diagonals of the red and blue rhombi of $S$, seen from above. See Figure 6 (f). Equivalently, they are the east and south pointing arrows in the six-vertex model of Figure 6 (c). They correspond to distinct adjacent entries in the CSM (see Figure 6 (g)) and to decreasing adjacent entries in the HFM (see Figure 6 (h)).
@@
  • The fully packed loop is given by the horizontal diagonals of the red and orange (resp. green and blue) rhombi of $S$ located at odd (resp. even) heights, seen from above. See Figure 6 (e).
@@
Figure 7 illustrates these correspondence on bigger examples of ASM (b), peak-pit model (e), fully packed loop (i), and cliff model (j).
alternating sign matrix peak-pit model Refer to caption Refer to caption Refer to caption Refer to caption fully packed loop cliff model
Figure 7. Bigger examples of the correspondence between ASM, peak-pit model, fully packed loop, and cliff model.
@@ 4.4. Catalan triangles
We note that our definition of Gelfand–Tsetlin patterns only differs from the classical definition by our use of barycentric coordinates to parametrize $\mathcal{D}_{n}$. We also observe that our two conditions for doubly gapless patterns are usually called gapless and left-gapless. Gapless triangles were introduced by A. Ayyer, R. Cori and D. Gouyou-Beauchamps [ACGB11], who described a bijection between gapless Gog triangles and gapless Magog triangles.
@@ 4.7. Symmetric and self-dual catalan triangles
$n$ 11 22 33 44 55 66 77 88 99 $\cdots$
$|\mathbf{Cat}_{n}|$ 11 22 66 $28$ $202$ $2252$ $38756$ $1028964$ $42127054$ $\cdots$
symmetric 11 22 44 $12$ $36$ $168$ $768$ $5556$ $38904$ $\cdots$
self-dual 11 0 22 0 $10$ 0 $120$ 0 $3434$ $\cdots$
@@ 6.1. Nappes
$n$ 11 22 33 44 55 66 77 88 99 $10$ $\cdots$
$N_{n}$ 0 11 99 $52$ $260$ $1291$ $6915$ $41814$ $289758$ $2291381$ $\cdots$
$O_{n}$ 0 11 88 $38$ $140$ $443$ $1268$ $3384$ $8584$ $20965$ $\cdots$
Table 3. First values of $N_{n}$ and $O_{n}$ from Proposition 6.7. Note that $O_{n}$ is [OEI10, A034009].
@@ 6.3. Minimal elements versus minimal permutations
$n$ 22 33 44 55 66 77 $\cdots$ all congruences 22 $16$ $1024$ $1048576$ $34359738368$ $72057594037927936$ $\cdots$ min permutational 22 $12$ $216$ $10480$ $1344096$ $465473984$ $\cdots$ min and max permutational 22 99 $69$ $716$ $8986$ $128065$ $\cdots$
Table 4. The numbers of congruences of $\mathbf{ASM}_{n}$ depending on whether we require all minima and/or maxima to be permutational. Note that the first line is [OEI10, A125791]
@@ Acknowledgments
We are grateful to Lucas Gagnon, Jean-Christophe Novelli, and Wenjie Fang for their insightful discussions and suggestions. We also thank Romain Di Vozzo for his helpful discussions and assistance with the 3D models.

### v2

@@ Abstract.
FH was partially supported by the European projet FRESCO (European Union’s Horizon 2020 research and innovation program grant No.101001995) and by the French–Austrian project PAGCAP (ANR-21-CE48-0020 & FWF I 5788). VP was partially supported by the Spanish project PID2022-137283NB-C21 of MCIN/AEI/10.13039/501100011033 / FEDER, UE, by the Spanish–German project COMPOTE (AEI PCI2024-155081-2 & DFG 541393733), by the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M), by the Departament de Recerca i Universitats de la Generalitat de Catalunya (2021 SGR 00697), and by the French–Austrian project PAGCAP (ANR-21-CE48-0020 & FWF I 5788).
@@ 1. Introduction
Our work starts from the observation that this excedance relation on the Bruhat order actually extends to a congruence of the lattice of alternating sign matrices [MRR82, MRR83, RR86] (which is the MacNeille completion of the Bruhat order [LS96]). The congruences of a lattice are the equivalence relations which respect the lattice structure, and they come with rich combinatorial and algebraic properties. Our main results (partially) extend the results of [BG24] to all lattice congruences $\equiv$ on alternating sign matrices whose quotient is isomorphic to the Stanley lattice, which we call catalan congruences.
@@
A few comments on this statement. In (i), the covexillary permutation matrices that appear as maxima of congruence classes depend on the congruence. However, any covexillary permutation matrix appears as the maxima of a class for at least one catalan congruence. In (ii), the minimal alternating sign matrix of each class is not necessarily a permutation (we actually discuss along the paper the specific congruences with this property). But the permutations in each congruence class form an interval of the Bruhat order, whose minimum is $321$-avoiding and whose maximum is covexillary, and (by counting) all $321$ permutations appear as minimal permutations of congruence classes. Finally, we note that the problem to extend Theorem 1.1 (iv) to any catalan congruence is under investigation.
Our approach to Theorem 1.2 is a combination of lattice and geometric perspectives on alternating sign matrices. It is well known that the lattice of alternating sign matrices is distributive, hence isomorphic to the inclusion lattice of lower sets of its join irreducible subposet by the fundamental theorem for distributive lattices. As already observed and exploited in [EKL+92, PRO01, STR11], the join irreducible poset of the lattice of alternating sign matrices can be seen geometrically as a poset on the integer points inside the (n2)(n-2)th dilate of a $3$-dimensional tetrahedron resting on an edge (see Figure 3). Pushing slightly further this geometric perspective in Section 2.3, we interpret a lower set in this tetrahedral poset as a stack of rhombic tetrahedra (see Figure 8). The upper hull of this stack is the colored polygonal surface separating the lower set from its complementary upper set (see Figure 5). Projecting this surface to the ground (see Figure 6) naturally leads to interpret geometrically the classical combinatorial models in bijection with alternating sign matrices (e.g. height functions, aztec diamond [EKL+92], six-vertex configurations [KUP96], osculating path configurations [BEH08], bumpless pipe dreams [WEI21], fully packed loops [BBN+96, DE 05]), and to cook up some seemingly new models (e.g. cliff configurations). We note that this perspective also naturally leads to higher dimensional analogues of alternating matrices which will deserve further investigation.
@@ 2.2. Dyck paths
A Dyck path of semilength $n$ is a path with up steps $(1,1)$ and down steps (1,1)(1,-1), starting at $(0,0)$, ending at $(2n,0)$, and never passing strictly below the horizontal axis. They define the Dyck path lattice $\mathbf{Dyck}_{n}$ (also known as the Stanley lattice) with the relation $P\leq Q$ if $P$ stays below $Q$. See Figure 2 (left). The join (resp. meet) irreducible Dyck paths are those with a single peak (resp. valley), hence the poset $\mathcal{J}(\mathbf{Dyck}_{n})$ (resp. $\mathcal{M}(\mathbf{Dyck}_{n})$) is isomorphic to the triangular poset $\mathcal{D}_{n}$ of triples $(y_{1},y_{2},y_{3})$ with $y_{1}+y_{2}+y_{3}=n-2$, ordered by $(y_{1},y_{2},y_{3})\leq(z_{1},z_{2},z_{3})$ if and only if $y_{1}+y_{2}\leq z_{1}+z_{2}$ and $y_{2}+y_{3}\leq z_{2}+z_{3}$. See Figure 3 (left). (Note that this poset could be encoded using two coordinates. However, we prefer to work with barycentric coordinates throughout the paper, as they make the presentation more symmetric.)
@@ 2.3. Alternating sign matrices
Note that via the above-mentioned bijection between ASMs and HFMs, we can transport the lattice $\mathbf{HFM}_{n}$ on HFMs of order $n$ to a lattice $\mathbf{ASM}_{n}$ on ASMs of order $n$. See Figure 2 (right). Identifying a permutation $\sigma$ of $[n]$ with its matrix $P_{\sigma}\mbox{\,\raisebox{0.86108pt}{\scriptsize$\mathrm{:}$}$=$\,}[\delta_{\sigma(i),j}]_{i,j\in[n]}$, this lattice $\mathbf{ASM}_{n}$ is the MacNeille completion of the Bruhat order on permutations of $[n]$ [LS96]. We call ascents and descents of an ASM the ascents and descents of its corresponding HFM. The join (resp. meet) irreducible ASMs are known to be the matrices of the bigrassmannian (resp. anti-bigrassmannian) permutations, i.e. the permutations $\sigma$ such that $\sigma$ and its inverse $\sigma^{-1}$ have only one descent (resp. ascent) [LS96]. Namely, an integer point $\boldsymbol{x}=(x_{1},x_{2},x_{3},x_{4})$ in the tetrahedron corresponds to the bigrassmannian and anti-bigrassmannian permutations
@@
It is also interesting to represent an ASM as the surface $S$ separating the lower set $X$ of $\mathcal{T}_{n}$ and its complement. For this, we include additional edges so that each integer point of the tetrahedron gets $12$ incident arrows (one in the direction $\boldsymbol{e}_{i}-\boldsymbol{e}_{j}$ for each $i\neq j\in[4]$). The horizontal edges, in direction $\pm(1,0,0,-1)$ and $\pm(0,1,-1,0)$, are colored in gray. The additional endpoints are considered in $X$ (resp. in its complement) if they are incident to an arrow which is also incident to one of the two bottom (resp. top) faces of the tetrahedron. We then add a rhombus orthogonal to each arrow from a point in $X$ to a point in its complement (the color of the rhombus matches the color of the orthogonal arrow). The corresponding surface is illustrated in Figure 5 (middle and right) [^1] See also https://www.ub.edu/comb/vincentpilaud/documents/publications/ASM.html..
@@
    Note that, up to recoloring, this rainbow tiling is equivalent to a 4-coloring of the aztec diamond [EKL+92].
@@
  • The six-vertex configuration [KUP96] is given by the horizontal diagonals of the rhombi of $S$, oriented counterclockwise around the surface $S$ (in other words, in the direction of the march of the dahu [^2] https://en.wikipedia.org/wiki/Dahu), and seen from above. See Figure 6 (c). Equivalently, it can be obtained by placing south/west/north/east pointing arrows in the red/blue/orange/green tiles of the rainbow tiling of Figure 6 (a). This gives an orientation of the integer grid such that
@@
    It has $6$ types of vertices [Uncaptioned image], hence the name. Note that, as the nonzero entries of the ASM correspond to saddle points of $S$, they correspond to the vertices of the six-vertex configuration with opposite incoming arrows and opposite outgoing arrows. See the connection between Figure 6 (b) & (c).
  • The six-square configuration is given by the non-horizontal diagonals of the rhombi of $S$, oriented upwards, and seen from above. See Figure 6 (d). Equivalently, it can be obtained by placing east/south/west/north pointing arrows in the red/blue/orange/green tiles of the rainbow tiling of Figure 6 (a). Equivalently, it is the dual directed graph of the six-vertex configuration of Figure 6 (c) (meaning that each arrow is rotated by $90^{\circ}$). This gives an orientation of the half-integer grid such that
@@
    It has $6$ types of squares [Uncaptioned image], hence the name.
  • The peak–pit matrix is given by the local optima of the surface $S$, seen from above. Figure 6 (e). They can be visualized as the rainbow half integer points of the rainbow tiling of Figure 6 (a). Equivalently, they are the points with indegree or outdegree $4$ in the six-square configuration of Figure 6 (d). We are not aware that this model has been studied before.
  • The osculating path configuration [BEH08] or bumpless pipe dream [WEI21] is given by the horizontal diagonals of the red and blue rhombi of $S$, seen from above. See Figure 6 (f). Equivalently, they are the east and south pointing arrows in the six-vertex configuration of Figure 6 (c). They correspond to distinct adjacent entries in the CSM (see Figure 6 (g)) and to decreasing adjacent entries in the HFM (see Figure 6 (h)).
@@
  • The fully packed loop [BBN+96, DE 05] is given by the horizontal diagonals of the red and orange (resp. green and blue) rhombi of $S$ located at odd (resp. even) heights, seen from above. See Figure 6 (e).
@@
Figure 7 illustrates these correspondence on bigger examples of ASM (b), peak-pit matrix (e), fully packed loop (i), and cliff configuration (j).
alternating sign matrix peak-pit matrix Refer to caption Refer to caption Refer to caption Refer to caption fully packed loop cliff configuration
Figure 7. Bigger examples of the correspondence between ASM, peak-pit matrix, fully packed loop, and cliff configuration.
@@ 4.4. Catalan triangles
We note that our definition of Gelfand–Tsetlin patterns only differs from the classical definition by our use of barycentric coordinates to parametrize $\mathcal{D}_{n}$. We also observe that our two conditions for doubly gapless patterns are usually called gapless and left-gapless. Gapless triangles were introduced by A. Ayyer, R. Cori and D. Gouyou-Beauchamps [ACG11], who described a bijection between gapless Gog triangles and gapless Magog triangles.
@@ 4.7. Symmetric and self-dual catalan triangles
$n$ 11 22 33 44 55 66 77 88 99 $\cdots$ OEIS
$|\mathbf{Cat}_{n}|$ 11 22 66 $28$ $202$ $2252$ $38756$ $1028964$ $42127054$ $\cdots$ A391969
symmetric 11 22 44 $12$ $36$ $168$ $768$ $5556$ $38904$ $\cdots$ A390493
self-dual 11 0 22 0 $10$ 0 $120$ 0 $3434$ $\cdots$ A391973
@@ 6.1. Nappes
$n$ 11 22 33 44 55 66 77 88 99 $10$ $\cdots$ OEIS
$N_{n}$ 0 11 99 $52$ $260$ $1291$ $6915$ $41814$ $289758$ $2291381$ $\cdots$ A391692
$O_{n}$ 0 11 88 $38$ $140$ $443$ $1268$ $3384$ $8584$ $20965$ $\cdots$ A034009
Table 3. First values of $N_{n}$ and $O_{n}$ from Proposition 6.7.
@@ 6.3. Minimal elements versus minimal permutations
$n$ 22 33 44 55 66 77 $\cdots$ OEIS all congruences 22 $16$ $1024$ $1048576$ $34359738368$ $72057594037927936$ $\cdots$ A125791 min permutational 22 $12$ $216$ $10480$ $1344096$ $465473984$ $\cdots$ A391690 min & max permutat. 22 99 $69$ $716$ $8986$ $128065$ $\cdots$ A391691
Table 4. The numbers of congruences of $\mathbf{ASM}_{n}$ depending on whether we require all minima and/or maxima to be permutational.
@@ Acknowledgments
We are grateful to Lucas Gagnon, Jean-Christophe Novelli, Wenjie Fang, and Jessica Striker for their insightful discussions and suggestions. We also thank Romain Di Vozzo and Selena Pere for their helpful discussions and assistance with the 3D models. We thank anonymous referees of the FPSAC’26 conference for comments on the presentation.